OpenStudy (anonymous):

Express cos6 theta in terms of sin theta i need a worked out example to solve this

OpenStudy (mayaal):

r u indian?

OpenStudy (anonymous):

Demoivers theorem and binomial theorem has to be used to solve this.. After simplifying to a step i couldn't simplify the i terms in the equation.. Please help me to solve this

OpenStudy (anonymous):

Yes I am..

OpenStudy (anonymous):

\[\cos^6\theta~\text{or}~\cos6\theta~?\]

OpenStudy (anonymous):

\[ \cos6 \theta \rightarrow \cos2 (3\theta)\] since \[\cos2\theta=1-2\sin^2\theta \rightarrow \cos2(3\theta)=1-2\sin^2(3\theta)\]

OpenStudy (anonymous):

It is cos 6 theta not to the power of 6 theta...